the sum of 2 numbers is 29 and their difference is 13. solve.
Answer by flame8855(424)

let x and y be the two numbers
x+y = 29
x-y = 13
solve simultaneously
x+y +x -y = 29 + 13
2x=42
x= 21
x+y = 29
y= 29-x= 29- 21 = 8
x is 21
y is 8